# Enhancing Multimodal Unified Representations for Cross Modal Generalization

### v2

@@ Abstract
To enhance the interpretability of multimodal unified representations, many studies have focused on discrete unified representations. These efforts typically start with contrastive learning and gradually extend to the disentanglement of modal information, achieving solid multimodal discrete unified representations. However, existing research often overlooks two critical issues: 1) The use of Euclidean distance for quantization in discrete representations often overlooks the important distinctions among different dimensions of features, resulting in redundant representations after quantization; 2) Different modalities have unique characteristics, and a uniform alignment approach does not fully exploit these traits. To address these issues, we propose Training-free Optimization of Codebook (TOC) and Fine and Coarse cross-modal Information Disentangling (FCID). These methods refine the unified discrete representations from pretraining and perform fine- and coarse-grained information disentanglement tailored to the specific characteristics of each modality, achieving significant performance improvements over previous state-of-the-art models.
@@
Hai Huang1thanks: Equal contribution., Yan Xia [^1] footnotemark: 1, Shengpeng Ji1, Shulei Wang1, Hanting Wang1, Minghui Fang1, Jieming Zhu2, Zhenhua Dong2, Sashuai Zhou1, Zhou Zhao1 1Zhejiang University, 2Huawei Noah’s Ark Lab Correspondence: haihuangcode@outlook.com
@@ 1 Introduction
To address the aforementioned issues, we propose TOC and FCID.
@@ 2 Related Work
Multi-Modal Unified Representation: Recent work on multi-modal unified representations includes approaches that align modalities into a shared latent space (Petridis et al., 2018; Sarkar and Etemad, 2022; Andonian et al., 2022) and train modal-general encoders for cross-modal extraction (Chen et al., 2020b; Wang et al., 2022). Cross-modal distillation enables knowledge transfer between modalities (Sarkar and Etemad, 2022; Pedersoli et al., 2022), while bridging techniques connect representation spaces for improved unified representations (Wang et al., 2023). To enhance interpretability, many works use codebooks or prototypes (Duan et al., 2022; Lu et al., 2022; Liu et al., 2021a; Zhao et al., 2022; Xia et al., 2024; Huang et al., 2024). For instance, Duan et al. (2022) applies Optimal Transport to map features to prototypes, Xia et al. (2024) maps multimodal sequences to a common discrete semantic space. Our FCID framework addresses the inherent differences between text and audio-visual modalities through decoupling, enhancing multimodal unified representations.
@@ 5.1.2 Downstream
The unified representation pre-trained models will be evaluated on several downstream tasks using different datasets. Cross-modal event classification on AVE dataset: (Tian et al., 2018) training on one modality and evaluating on another. Cross-modal event localization on AVVP dataset: (Tian et al., 2020) localizing events in one modality and transferring to the other. Cross-dataset localization/classification: training on classification in AVE and evaluating localization in AVVP, transferring across datasets. Cross-modal classification between UCF-101 (Soomro et al., 2012) visual clips and VGGSound (Chen et al., 2020a) audio clips. The decoder in all of the above experiments consists of a single linear layer. Cross-modal Zero-shot Retrieval: We adopt a process similar to the test set (Yu et al., 2018), which consists of 500 pairs from MSCOCO (Chen and Dolan, 2011), assessing zero-shot retrieval capability for visual-text alignment. Clotho (Drossos et al., 2020) assesses zero-shot retrieval capability for audio-text alignment. Flickr Sound (Senocak et al., 2018) assesses zero-shot retrieval capability for audio-visual alignment. Cross-modal Generation: We use a 2-layer MLP and the IP-Adapter (Ye et al., 2023) as downstream decoders. The MLP is trainable during training but frozen during testing, while the IP-Adapter remains frozen throughout. By leveraging IP-Adapter’s image-to-image capability, the MLP bridges multimodal unified representation and image generation, enabling audio-to-image and text-to-image generation during testing. The model was fine-tuned on 4,500 FlickrSound (Senocak et al., 2018) image-audio pairs over 80,000 steps with a batch size of 8, and evaluated on 500 additional pairs. For details on the compared works, evaluation metrics, and hyperparameter settings, please refer to Appendix A.
@@ B.2 Computational efficiency
DCID 7.90GB 1.72h 5
CMCM 6.25GB 1.41h 8
FCID 7.42GB 1.74h 5
Table 8: Model Comparison: GPU Memory Usage, Time per Epoch, and Total Epochs
@@ B.6 Generation
Figure 11: More results of Cross-modal generation.

### v3

@@ Abstract
To enhance the interpretability of multimodal unified representations, many studies have focused on discrete unified representations. These efforts typically start with contrastive learning and gradually extend to the disentanglement of modal information, achieving solid multimodal discrete unified representations. However, existing research often overlooks two critical issues: 1) The use of Euclidean distance for quantization in discrete representations often overlooks the important distinctions among different dimensions of features, resulting in redundant representations after quantization; 2) Different modalities have unique characteristics, and a uniform alignment approach does not fully exploit these traits. To address these issues, we propose Training-free Optimization of Codebook (TOC) and Fine and Coarse cross-modal Information Disentangling (FCID). These methods refine the unified discrete representations from pretraining and perform fine- and coarse-grained information disentanglement tailored to the specific characteristics of each modality, achieving significant performance improvements over previous state-of-the-art models. The code is available at https://github.com/haihuangcode/CMG.
@@
Hai Huang1thanks: Equal Contribution, Yan Xia [^1] footnotemark: 1, Shengpeng Ji1, Shulei Wang1, Hanting Wang1, Minghui Fang1, Jieming Zhu2thanks: Project Lead, Zhenhua Dong2, Sashuai Zhou1, Zhou Zhao1thanks: Corresponding Author 1Zhejiang University,  2Huawei Noah’s Ark Lab haihuangcode@outlook.com
@@ 1 Introduction
To address the aforementioned issues, we propose two techniques: TOC and FCID.
@@ 2 Related Work
Multi-Modal Unified Representation: Recent work on multi-modal unified representations includes approaches that align modalities into a shared latent space (Petridis et al., 2018; Sarkar and Etemad, 2022; Andonian et al., 2022) and train modal-general encoders for cross-modal extraction (Chen et al., 2020b; Wang et al., 2022). Cross-modal distillation enables knowledge transfer between modalities (Sarkar and Etemad, 2022; Pedersoli et al., 2022), while bridging techniques connect representation spaces for improved unified representations (Wang et al., 2023). To enhance interpretability, many works use codebooks or prototypes (Duan et al., 2022; Lu et al., 2022; Liu et al., 2021a; Zhao et al., 2022; Xia et al., 2024; Huang et al., 2024; Fang et al., 2024). For instance, Duan et al. (2022) applies Optimal Transport to map features to prototypes, Xia et al. (2024) maps multimodal sequences to a common discrete semantic space. Our FCID framework addresses the inherent differences between text and audio-visual modalities through decoupling, enhancing multimodal unified representations.
@@ 5.1.2 Downstream
The unified representation pre-trained models will be evaluated on several downstream tasks using different datasets. Cross-modal event classification on AVE dataset: (Tian et al., 2018) training on one modality and evaluating on another. Cross-modal event localization on AVVP dataset: (Tian et al., 2020) localizing events in one modality and transferring to the other. Cross-dataset localization/classification: training on classification in AVE and evaluating localization in AVVP, transferring across datasets. Cross-modal classification between UCF-101 (Soomro et al., 2012) visual clips and VGGSound (Chen et al., 2020a) audio clips. The decoder in all of the above experiments consists of a single linear layer. Cross-modal Zero-shot Retrieval: We adopt a process similar to the test set (Yu et al., 2018), which consists of 500 pairs from MSCOCO (Chen and Dolan, 2011), assessing zero-shot retrieval capability for visual-text alignment. Clotho (Drossos et al., 2020) assesses zero-shot retrieval capability for audio-text alignment. Flickr Sound (Senocak et al., 2018) assesses zero-shot retrieval capability for audio-visual alignment. Cross-modal Generation: We use a 2-layer MLP and the IP-Adapter (Ye et al., 2023) as downstream decoders. The MLP is trainable during training but frozen during testing, while the IP-Adapter remains frozen throughout. By leveraging IP-Adapter’s image-to-image capability, the MLP bridges multimodal unified representation and image generation, enabling audio-to-image and text-to-image generation during testing. The model was fine-tuned on 4,500 FlickrSound (Senocak et al., 2018) image-audio pairs over 80,000 steps with a batch size of 8, and evaluated on 500 additional pairs. Please refer to Appendix A for details on the compared works, evaluation metrics, and hyperparameter settings.
@@ 6 Conclusion
Acknowledgments
We thank MindSpore (http://mindspore.cn) for the partial support of this work, which is a new deep learning computing framework.
@@ B.2 Computational efficiency
DCID 100% 100% 5
CMCM 81% 76% 8
FCID 96% 99% 5
Table 8: Model Comparison: Relative GPU Memory Usage and Time per Epoch (Compared to DCID), and Total Training Epochs
@@ B.5 Reconstruction
Figure 11: More results of cross-modal generation.